# VIN-NBV: A View Introspection Network for Next-Best-View Selection for Resource-Efficient 3D Reconstruction

### v1

@@ 4.3 Generalization Across Object Categories
In Fig.6 we see that when we evaluate our VIN-NBV policy on additional object categories, our policy is able to consistently outperform the coverage baseline. Larger gaps in performance occur early on, with both methods reaching similar final chamfer distance values at 20 captures. Our policy does particularly well, compared to the coverage baseline, for objects with more complex self-occlusions such as the toy motorcycle class. We include visual results of our policy and the coverage baseline after 10 acquisitions on several objects and from various viewing angles in Fig.LABEL:fig:10_acquisition_results_page_3, Fig.LABEL:fig:10_acquisition_results_page_1, Fig.LABEL:fig:10_acquisition_results_page_2, Fig.LABEL:fig:10_acquisition_results_page_4.
(a) Dinosaurs
(b) Toy Motorcycles
(c) Toy Animals
Figure 6: We graph the average chamfer distance at each capture stage in three additional object classes namely dinosaurs, toy motorcycles, and toy animals. We compare VIN-NBV to a coverage baseline for 20 captures.
(a) VIN Score predictions
(b) Oracle-RRI score predictions
Figure 7: 3D heatmap (top-down view) indicating the fitness score predicted by (a) our VIN policy and (b) the Oracle-NBV during the first capture step on house 1 from OmniObject3D [43]. Each point marks a query-view camera position, colored by its predicted reconstruction improvement score (blue=low, red=high). While VIN-NBV has high precision, accurately identifying the most helpful viewpoints, it has low recall, often missing many useful views.
@@ 5 Limitations
RGB-D vs RGB capture. We use ground truth depth maps in RGB-D capture to match the evaluation setting of the prior works GenNBV [25] and ScanRL [26], but these depth maps are noise-free and not reflective of true depth sensor readings. We observe that depth maps estimated from RGB images using Monocular Depth Estimation [46, 47] often lack multiview consistency and struggle with handling the scale of different objects especially with our synthetic data. Depth maps from MultiView Stereo algorithms [48, 49] often require more than 3 views and perform poorly with just 2 base images. Since VIN relies on 3D-aware featurization, inaccurate and multiview inconsistent depth maps can lead to a poorly reconstructed 3D scene and featurization, which the VIN network struggles to handle.
Performance gap with Oracle-NBV. Our method shows a significant gap to our Oracle-NBV baseline in the early capture stages, indicating that there is still room for improvement. In Fig. 4(a) we see that the Oracle baseline is able to make significantly better choices early on and converge much faster than our policy. In Fig. 4(b), we see that even with less time, the oracle baseline is still able to outperform our policy, and as more time is added, it converges much sooner. These gaps may indicate that the VIN is unable to properly mimic the behavior of the Oracle-NBV.
To understand this, in Fig. 7, we make a simple visualization of the 3D heatmap of Relative Reconstruction Improvement scores of densely sampled query views around the object, calculated by the VIN-NBV and the Oracle-NBV, for the first capture step of house 1 from OmniObject3D. We observe that while Oracle-NBV calculates a relatively large set of query views to be effective with high fitness scores (marked in red), VIN-NBV only predicts a subset of these views to be effective, and misses out on a large set of potentially good views. Hence, during evaluation, the chances of VIN-NBV finding a ‘good’ view from a small set of sampled views is low, and can explain the performance gap between Oracle-NBV and VIN-NBV.
@@ 6 Conclusion
In this paper, we revisited the Next Best View problem and proposed the VIN and VIN-NBV policy, a generalizable policy that can determine a set of optimal acquisitions to maximize reconstruction quality. Our policy can be easily adapted to operate with limitations on the number of acquisitions or a time limit on robot motion. By optimizing for reconstruction quality rather than traditional coverage, VIN-NBV improves final reconstruction results without requiring prior scene knowledge, extra image captures, per-scene training, or complex RL policies. Evaluations on OmniObject3D [43] show VIN-NBV outperforms state-of-the-art RL methods, reducing reconstruction error by up to 40%. VIN-NBV can enable robots to efficiently acquire necessary 3D information for various applications where time is of the essence.
See pages 3 of Figures/extra_visuals.pdf
See pages 2 of Figures/extra_visuals.pdf
See pages 1 of Figures/extra_visuals.pdf
See pages 4 of Figures/extra_visuals.pdf

### v3

@@ 4.4 Generalization Across Object Categories
In Fig.8 we see that when we evaluate our VIN-NBV policy on additional object categories, it is able to consistently outperform the coverage baseline. Larger gaps in performance occur early on, with both methods reaching similar final Chamfer Distance values at 20 captures. Our policy does particularly well, compared to the coverage baseline, for objects with more complex self-occlusions such as toy motorcycles. We include visual results of our policy and the coverage baseline after 10 acquisitions on several objects and from various viewing angles in Appendix 7.3.
Policy Dinosaurs Trucks Animals
Scan-RL [29] 0.06 0.88 0.17
GenNBV [5] 0.03 0.81 0.16
VIN-NBV (ours) 0.04 0.57 0.08
Table 2: Reconstruction quality comparisons to prior works, with Chamfer Distances ($\downarrow$ is better), on 3 additional classes from OmniObject3D [41] after 20 total acquisitions.
The toy truck category was also evaluated for final Chamfer Distance so that direct comparisons against GenNBV’s [5] reported results could be made in Table 2. VIN-NBV outperforms ScanRL but is slightly behind GenNBV on dinosaurs, while significantly beating both baselines on toy animals and trucks.
@@ 5 Limitations
We use ground truth depth maps in RGB-D capture to match the evaluation setting of the prior works GenNBV [5] and ScanRL [29], but these depth maps are noise-free and not reflective of true depth sensor readings. The performance of our method has not been evaluated with real world data and could have different performance in such settings. In Fig. 4 we see a gap to the Oracle-NBV in early capture stages, indicating that our method still has room for improvement.
@@ 6 Conclusion
In this paper, we revisit the Next Best View problem and propose the VIN and VIN-NBV policy, a generalizable policy that can determine a set of optimal acquisitions to maximize reconstruction quality. Our policy can be easily adapted to operate with limitations on the number of acquisitions or a time limit on agent motion. By optimizing for reconstruction quality rather than traditional coverage, the VIN-NBV policy improves final reconstruction results without requiring prior scene knowledge, extra image captures, per-scene training, or complex RL policies. Evaluations on OmniObject3D [41] show VIN-NBV outperforms state-of-the-art RL coverage based methods, reducing reconstruction error by up to 40%. The VIN-NBV policy can enable agents to efficiently acquire necessary 3D information for various applications where time is of the essence.